OpenStudy (anonymous):

the circle C has eqn (x-3)^2+(y-4)^2=25 and the eqn of tangent at the point (6,8) is y=4/3 x ......calculate the co-ordinates of the point of intersection of C with the straight line y=2x+3

OpenStudy (anonymous):

put y=2x +3 in the equation of circle to get the values of y

OpenStudy (anonymous):

(x^2+9 - 6x) +(2x-1)^2 =25

OpenStudy (anonymous):

x= -1, 3 on solving

OpenStudy (anonymous):

now put these values in y

OpenStudy (anonymous):

fan me pls if u like answer

OpenStudy (anonymous):

already fanned u

OpenStudy (anonymous):

thanx

OpenStudy (anonymous):

the points P and Q have position vectors relative to the origin O given respectively by p=-i+6j and q=3i+8j.....a)calculate in degrees the size of the acute angle between p and q...b)calculate the area of triangle POQ